# An Evaluation of N-Gram Selection Strategies for Regular Expression Indexing in Contemporary Text Analysis Tasks [Experiment, Analysis & Benchmark]

### v1

@@ 1. Introduction
Existing solutions, such as frequency-based selection (prioritizing common n-grams) (Cho and Rajagopalan, 2002), coverage-optimized methods (minimizing redundant matches) (Hore et al., 2004), and heuristic-driven strategies like LPMS (Tsang and Chawla, 2011), were proposed decades ago. For instance, the FREE (Cho and Rajagopalan, 2002) and BEST (Hore et al., 2004) methods, introduced in the early 2000s, focus on minimizing computational space overhead and index size for memory-constrained systems. Despite being an active area of research, surprisingly, the performance of these methods on modern hardware is lacking for three reasons. The absence of a comprehensive, unbiased, and systematic understanding of these methods is due to three main reasons. First, existing methods have not been evaluated within the same experimental settings (such as the same hardware or using the same B+-Tree implementation), leading to incomplete comparisons. Second, prior work has only compared specific metrics and on a limited number of datasets. Third, the absence of a standardized and inclusive testing framework hinders the investigation and testing of methods in practical implementations. Consequently, the lack of a modern comparison of these past methods has led to the adoption of these legacy strategies in recent studies (Qiu et al., 2022; Chockchowwat et al., 2022) without re-evaluating their assumptions. These older methods may not be appropriate when dealing with larger datasets (now common in applications domains such as genenomics, IoT/sensor, and log analytics) as it stresses their n-gram selection methods. For example, as we will see later, BEST’s quadratic runtime complexity becomes prohibitive on even a small 10GB text corpus, forcing practitioners to downsample data or abandon indexing altogether. In fact, n-gram indexing used by Postgres (Christopher Kings-Lynne and Korotkov, [n.d.]) and GitLab code search (Peterse, 2016) abandon the n-gram selection strategies and build the index with all n-grams for only trigrams (i.e. $n=3$) or for several fixed values $n$ (for instance, indexing all bigrams and trigrams). In summary, there are three key limitations that motivate our work.
@@
    Lack of Comprehensive Comparison. Prior evaluations focus on specific metrics (e.g., false-positive rate) or evaluate their method using only a single workload, failing to capture the diverse demands of real-world workloads. For example, genomic databases prioritize minimizing false positives due to a large number of repeated queries on the same dataset, while log analysis prioritizes rapid index rebuild overhead for large scale streaming data analysis. No study has systematically compared and contrasted the three key methods FREE, BEST, and LPMS across wide-ranging scenarios.
@@
2. Methods and Benchmark Framework Implementation. In the absence of original code, we implemented FREE, BEST, and LPMS as described in the respective papers, using modern C++. While BEST’s original design includes parallelism for multi-core CPUs, FREE and LPMS lack scalable implementations. We modernized all three methods by: (1) redesigning FREE with parallelism to leverage modern multi-core CPUs, and (2) implementing LPMS with Gurobi, which inherently supports multi-threading in its LP solver. Additionally, we introduced an early stopping mechanism to all three methods to configure the maximum number of n-grams selected. We also developed an end-to-end benchmark framework that manages data loading, method selection, configuration settings, regex matching, and result reporting. This framework is modular, easily extensible, and publicly available at the link noted in the abstract. All experiments were conducted on a multi-core machine with large memory, addressing (M.2).
3. Broad Workload Benchmark. To address the lack of comprehensive empirical comparisons on workloads with different characteristics, we evaluate the methods across five workloads. This includes using three legacy benchmarks from prior work and two real-world workloads in contemporary regex matching tasks.
@@ 2. Related Work
N-gram indexing was introduced to further improve regex query performance by precomputing substrings of length n as index keys. Besides being used for pattern matching, n-gram indexing is also common in approximate string searching. Early works compared different n values and decided to index all trigrams. Other works index n-grams with different n values for various workloads. The database community has conducted extensive research on effective indexing strategies for regular expression queries. While existing works primarily focus on improving regex query performance through optimized index structures, relatively fewer studies have explored n-gram key selection strategies for regex indexing. N-gram indexing has also been used for the closely related problem of indexing regular expressions (instead of data) to find out which regexes match a given input string (Chan et al., 2003). Similar ideas have also been used for indexing data to speed up regular path query evaluation (Li et al., 2001).
The first known works discussing n-gram selection for regex indexing is FREE (Cho and Rajagopalan, 2002). The high-level idea of selecting a covering n-gram set was proposed in a study of Asian language indexing (Ogawa and Matsuda, 1998). The concept of a covering set was further developed by Kim et al. (Kim et al., 2010) for a more efficient n-gram selection method considering I/O cost. The analogy between the set-covering problem and the n-gram selection problem was formalized in BEST (Hore et al., 2004), where the authors presented a near-optimal algorithm to select variable-length n-grams considering index size constraints. Later work, LPMS (Tsang and Chawla, 2011), combined the solution formulations of FREE and BEST and reformulated the n-gram selection problem into linear programs. In this work, we compare the n-gram selection and indexing methods of FREE, BEST, and LPMS for in-memory workloads and indices.
@@ 3. Problem Definition
Definition 3.1.
N-Gram For a given finite alphabet $\Sigma$, an n-gram $g$ is a sequence of $n$ characters $g=g_{1}g_{2}\cdots g_{n}$ where $g_{i}\in\Sigma$.
We use $|g|$ to denote the length of the n-gram $g$. A literal is a string from the set $\Sigma*$, where * is the standard closure operator. A common operation on a regex query $q$ is to identify the maximal literal components from the regex. For a literal $l$, we will use $G(l)$ to denote the set of all possible substrings of $l$. The set of all possible n-grams of a regex $q$ is $G(r)=\bigcup_{\text{maximal }l\in q}G(l)$.
@@
A workload $W=(Q,D)$ consists of the set of regex query $Q=\{q_{1},q_{2},\cdots\}$ and dataset $D=\{d_{1},d_{2},\cdots\},d_{i}\in\Sigma*$. The size of the dataset is defined as $|D|=\sum_{d_{i}\in D}|d_{i}|$. Each $d_{i}$ will be referred to as a data record. Next, we define the support of an n-gram.
Definition 3.3.
Support The support $s$ of a literal $g$ in dataset $D$ is the number of elements in $D$ that contains $g$. Similarly, its support in query set $Q$ is the number of individual queries which contains $g$ as a literal.
@@
Definition 3.4.
Selectivity The selectivity $c$ of an n-gram $g$ in string set $D$ is the fraction of individual strings in $D$ that contains $g$.
$c_{g}=\frac{s_{D}(g)}{|D|}$
@@ 4. Methods Overview
Method N-Gram Source Selection Criteria Selectivity Threshold ($c$) N-Gram Constraint Index Structure Index Size Constraint
FREE (Cho and Rajagopalan, 2002) dataset prefix-free, selectivity 0.1 $2\leq n\leq 10$ Inverted Index -
BEST (Hore et al., 2004) queries & dataset Utility(g) = benefit($g$)/cost($g$) $0.05\leq c\leq 0.1$ B+-Tree Index User defined
LPMS (Tsang and Chawla, 2011) queries & dataset prefix-free and utility optimized - - Inverted Index -
@@ 4.1.1. Selection Strategy
FREE uses the dataset in the workload as the source of n-gram selection, and selects a prefix-free set of n-grams based on selectivity. Note that FREE does not use the query workload for n-gram selection. The candidate set $G(W)$ of n-grams is $G(W)=\bigcup_{d\in D}G(d)$. For the candidates $G(W)$, FREE decides whether a n-gram will be selected by Usefulness.
Definition 4.1 (FREE).
Usefulness By setting a fixed selectivity threshold $\mathbf{c}$, n-gram $g$ is deemed useful if its selectivity is less than the selectivity threshold, $c_{g}<\mathbf{c}$.
FREE makes two important assumptions when selecting n-grams for indexing.
@@
FREE selects only n-grams that are useful. Since usefulness selection criteria is essentially a selectivity upper bound, if we know that a n-gram $g$ is useful, then any longer n-gram having $g$ as a substring is also useful. For example, if n-gram pdf is useful with a selectivity $c_{g}$, then n-grams such as $g^{\prime}=$pdf and $g^{\prime\prime}=$ pdfg will have selectivity $0\leq c(g^{\prime})<\mathbf{c}$ and $0\leq c(g^{\prime\prime})<\mathbf{c}$, and therefore also useful.
@@
For an n-gram $g=g_{1}g_{2}\cdots g_{n_{1}}$ of size $n_{1}$ having selectivity $c_{g}$, any other longer n-grams of size $n_{2}$
@@
To effectively reduce the size of n-grams selected, FREE selects the prefix-minimal useful set of n-grams, deriving from the Apriori method of finding the maximal frequent sets in data mining literature (Agrawal et al., 1996). Essentially, it generates all candidate n-grams in increasing size order iteratively. For iteration $i$, it generates the candidate set of n-grams of length $i$ by extending all useless n-grams of length $i-1$ by one character, inserting the useful n-grams in the candidate set into the index, and using the useless ones in the candidate set for iteration $i+1$. This way, it is not necessary to generate all possible n-grams in each iteration. This method also ensures that the set of n-grams in the index is a prefix-minimal set, as the breadth-first search ensures that the shortest prefix of a useful n-gram is visited first. No n-grams in candidate sets of future iterations have the selected n-grams as prefixes, as the useful n-grams are never extended to generate a candidate n-gram.
FREE confines the search space by constraining the length $n$ of the n-grams selected. Briefly, there are two parameters to tune: 1) selectivity threshold $c$ that distinguish useful n-grams from useless n-grams, and 2) n-gram size $n$ to control the length of the index keys. In the original paper, the authors uses $c=0.1$ and $2\leq n\leq 10$ in their experiments. The accompanied index data structure used is inverted index with n-grams as keys and posting lists as values.
(a) Literal based query plan before checking index keys.
(b) Final query plan after index scan.
Figure 1. Example index search plan tree for regex <ahref=("|’).*ZZZ\. pdf("|’)> built by FREE, where the n-grams ", , Z, and pdf are indexed.
4.1.2. Regex Compilation
We provide an example index search query plan in Figure 1 for the example regex, <ahref=("|’).*ZZZ\. pdf("|’)>, assuming that only ", , Z, and pdf are indexed. The literal tree shown in Figure 1(a) is built sorely by extracting all literals from the regex. Referring to the index, we notice that <ahref= and > are not indexed, and two n-grams in the literal ZZZ.pdf, Z and pdf are indexed. We simplify the plan to get Figure 1(b).
@@ 4.1.3. Complexity Analysis
Let us denote the candidate set of each step $i$ as $M_{i}$. In the initial step, FREE scans the entire dataset to find all unique uni-grams and calculates their selectivities. The space needed is a temporary hashmap with the key being all unique unigrams of size $|M_{1}|=|\Sigma|$. The time complexity for the base step is $O(|D|+|\Sigma|)$. After scanning through the dataset, a pass is made on all unique unigrams to calculate their selectivity and move the useful unigrams from the candidate set into the index. In the next iteration, the remaining useless uni-grams in the candidate set of the previous step are extended by one character to construct a candidate set of bi-grams, $M_{2}$. The runtime space overhead of this iteration is $O(|M_{2}|)$, and the compute time overhead is $O(|D|+|M_{2}|)$.
For each iteration $i$, we have the runtime space overhead $O(|M_{i}|)$ and compute overhead $O(|D|+|M_{i}|)$. Since each step $i$ selects all useful n-grams of size $i$, and the temporary hashmaps of previous iterations are never revisited after candidate set $i$ is generated, for a FREE n-gram selection with $n\leq k$ for some integer $k$, we have the runtime space overhead $O(|M_{k}|)$ and compute overhead $O(k\cdot(|D|+|M_{k}|))$ for each iteration. As the number of n-grams of any size on a dataset is at most the number of characters in the dataset, we have the overall runtime space overhead $O(|D|)$ and the compute overhead as $O(2\cdot|D|)$ for all iterations.
@@ 4.2.1. Selection Strategy
Definition 4.3 (BEST).
Cover For a workload $W=(Q,D)$, the cover of an n-gram $g$ is the set of data records $d\in D$ and query $q\in Q$ pairs defined as follows:
$cover(g)=\{(q,d)\quad|\quad g\in q\land g\notin d,\quad\forall(q,d)\in Q\times D\}$
For a set of n-grams $G$, the cover of the n-gram set is the union of the cover of each n-gram in the set.
$cover(G)=\bigcup_{g\in G}cover(g)$
The n-gram selection problem is then transformed to a budgeted maximum set cover problem (Khuller et al., 1999) that aims to find the minimum set $G$ that achieves maximum cover. Formally, BEST define the value of indexing an n-gram with respective to the entire workload by its benefit.
Definition 4.4 (BEST).
Benefit For a workload $W=(Q,D)$ and an n-gram set $I$, the benefit of an n-gram $g\notin I$ is the number of additional query-data pairs covered by $g$ that is not already covered by $I$.
@@
Definition 4.5 (BEST).
Cost For a workload $W$ consists of dataset $D$ and query set $Q$, the cost to index an n-gram $g$ is the storage overhead of $g$.
@@
Definition 4.6 (BEST).
Utility For a workload $W$ consists of dataset $D$ and query set $Q$ and an existing n-gram set $I$ on $W$, the utility of indexing an additional n-gram $g$ is the ratio of its benefit over cost.
$utility(g)=\frac{benefit(g,I)}{cost(g)}$
BEST selects n-grams based on their utility as defined above. The brute-force method would be to iteratively select the n-gram with the highest utility, $g_{max}$, based on the workload $W$ and the set of already-selected n-grams $I$, among all possible n-grams with a positive benefit.
@@
Therefore, according to the sparsity assumption Assumption BEST-3, BEST choose to use adjacency lists: Q-G-list and G-D-list rather than two matrices of sizes $|Q|\cdot|G|$ and $|G|\cdot|D|$ to store the existence of n-grams in the workload to reduce space usage. The candidate n-grams, queries, and data records are each assigned an index number. The indices of Q-G-list correspond to query numbers, and each element is a list of n-gram numbers in this query. Similarly, the indices of G-D-list correspond to n-gram numbers, and each element is a list of data record numbers which the n-gram is in.

### v2

# An Evaluation of N-Gram Selection Strategies for Regular Expression Indexing in Contemporary Text Analysis Tasks. Extended Version

@@ 1. Introduction
Legacy methods like frequency-based selection (Cho and Rajagopalan, 2002), coverage-optimized strategies (Hore et al., 2004), and heuristics like LPMS (Tsang and Chawla, 2011) were introduced decades ago to reduce index size and memory usage (Cho and Rajagopalan, 2002; Hore et al., 2004). Despite being an active area of research, surprisingly, the performance of these methods on modern hardware is lacking for three reasons. The absence of a comprehensive, unbiased, and systematic understanding of these methods is due to three main reasons.: (1) inconsistent experimental setups hinder fair comparisons, (2) evaluations are narrow in scope and datasets, and (3) a lack of standardized testing framework hindering practical assessment. As a result, recent studies (Qiu et al., 2022; Chockchowwat et al., 2022) continue to adopt these outdated approaches without re-evaluating their assumptions. These methods struggle with today’s large datasets (e.g., genomics, IoT, logs), where selection strategies break down. For instance, BEST’s quadratic runtime becomes impractical on even 10GB of text, often forcing downsampling or skipping indexing. Modern systems like Postgres (Christopher Kings-Lynne and Korotkov, [n.d.]) and GitLab (Peterse, 2016) bypass selection entirely, indexing all trigrams or fixed n-grams. These limitations motivate our work.
@@
    Lack of Comprehensive Comparison. Prior evaluations focus on specific metrics (e.g., false-positive rate) or evaluate their method using only a single workload, failing to capture the diverse demands of real-world workloads. No study has systematically compared and contrasted the three key methods FREE, BEST, and LPMS across wide-ranging scenarios.
@@
2. Methods and Benchmark Framework Implementation. In the absence of original code, we implemented FREE, BEST, and LPMS as described in the respective papers, using modern C++. While BEST’s original design includes parallelism for multi-core CPUs, FREE and LPMS lack scalable implementations. We modernized all three methods by: (1) redesigning FREE with parallelism to leverage modern multi-core CPUs, and (2) implementing LPMS with Gurobi, which inherently supports multi-threading in its LP solver. Our benchmarking framework is modular, easily extensible, and publicly available at the link noted in the abstract. All experiments were conducted on a multi-core machine with large memory, addressing (M.2).
3. Broad Workload Benchmark. To address the lack of comprehensive empirical comparisons on workloads with different characteristics, we evaluate the methods across eight workloads. This includes using three legacy benchmarks from prior work and two real-world workloads in contemporary regex matching tasks.
@@ 2. Related Work
N-gram indexing was introduced to further improve regex query performance by precomputing substrings of length n as index keys. Besides being used for pattern matching, n-gram indexing is also common in approximate string searching. Early works compared different n values and decided to index all trigrams. We refer to this method as Trigrams, and will use it for comparison in Section 6. Other works index n-grams with different n values for various workloads. The database community has conducted extensive research on effective indexing strategies for regular expression queries. While existing works primarily focus on improving regex query performance through optimized index structures, relatively fewer studies have explored n-gram key selection strategies for regex indexing. N-gram indexing has also been used for the closely related problem of indexing regular expressions (instead of data) to find out which regexes match a given input string (Chan et al., 2003). Similar ideas have also been used for indexing data to speed up regular path query evaluation (Li et al., 2001).
The first known work discussing n-gram selection for regex indexing is FREE (Cho and Rajagopalan, 2002). The high-level idea of selecting a covering n-gram set was proposed in a study of Asian language indexing (Ogawa and Matsuda, 1998). The concept of a covering set was further developed by Kim et al. (Kim et al., 2010) for a more efficient n-gram selection method considering I/O cost. The analogy between the set-covering problem and the n-gram selection problem was formalized in BEST (Hore et al., 2004), where the authors presented a near-optimal algorithm to select variable-length n-grams considering index size constraints. Later work, LPMS (Tsang and Chawla, 2011), combined the solution formulations of FREE and BEST and reformulated the n-gram selection problem into linear programs. In this work, we compare the n-gram selection and indexing methods of FREE, BEST, and LPMS for in-memory workloads and indices.
Several works have explored ML-based approaches for regex evaluation in security contexts. On resource-constrained edge devices, compressed on-device ML models are commonly used for high-throughput malicious packet detection. However, these methods rely on extensive preprocessing and assume access to high-quality training data—assumptions that don’t hold in log analytics, where analysis is routinely ad hoc. For a broader overview, see the survey by Xu et al. (Xu et al., 2016).
@@ 3. Problem Definition
Definition 3.1 (N-Gram).
For a given finite alphabet $\Sigma$, an n-gram $g$ is a sequence of $n$ characters $g=g_{1}g_{2}\cdots g_{n}$ where $g_{i}\in\Sigma$.
We use $|g|$ to denote the length of the n-gram $g$. A literal is a string from the set $\Sigma*$, where * is the standard closure operator. A common operation on a regex query $q$ is to identify the maximal literal components from the regex. For a literal $l$, we will use $G(l)$ to denote the set of all possible substrings of $l$. The set of all possible n-grams of a regex $q$ is $G(q)=\bigcup_{\text{maximal }l\in q}G(l)$. We use $G$ to define the set of all candidate n-grams considered in an n-gram selection method. Note that $G$ may be different for different methods, and we will further discuss it in Section 4. $G_{i}$ to is used to denote subset of candidate n-grams that has length exactly $i$.
@@
A workload $W=(Q,D)$ consists of the set of regex queries $Q=\{q_{1},q_{2},\cdots\}$ and dataset $D=\{d_{1},d_{2},\cdots\},d_{i}\in\Sigma*$. The size of the dataset is defined as $|D|=\sum_{d_{i}\in D}|d_{i}|$. Each $d_{i}$ will be referred to as a data record. Next, we define the support of an n-gram.
Definition 3.3 (Support).
The support $s$ of a literal $g$ in dataset $D$ is the number of elements in $D$ that contains $g$. Similarly, its support in query set $Q$ is the number of individual queries which contains $g$ as a literal.
@@
Definition 3.4 (Selectivity).
The selectivity $c$ of an n-gram $g$ in string set $D$ is the fraction of individual strings in $D$ that contains $g$, i.e., $c_{g}={s_{D}(g)}/{|D|}$.
@@ 4. Methods Overview
Method Source Selection Criteria Selectivity Threshold ($c$) N-Gram Constraint Time Complexity Space Complexity
FREE (Cho and Rajagopalan, 2002) $D$ Prefix-free Selectivity 0.1 $[2,10]$ $O(k\cdot|D|)$ $O(|D|)$
BEST (Hore et al., 2004) $Q,D$ Utility $[0.05,0.1]$ - $O(|D|+|Q|)$ $O(|I|\cdot|G(Q)|\cdot|D|\cdot|Q|)$
LPMS (Tsang and Chawla, 2011) $Q,D$ Prefix-free Utility optimized - - $O(\sum_{i\in[10]}\{|G_{i}(Q)|^{2.5}$ $\quad+|G_{i}(Q)|\cdot(|D|+|Q|)\})$ $O(\max_{i\in[10]}|G_{i}(Q)|\cdot(|D|+|Q|)$
VGGraph (Qiu et al., 2025) $Q,D$ Combines FREE and BEST - $2,\inf)$ $O(n\cdot|D|+|Q|\cdot|G(Q)|\cdot n)$ $O(|D|+|G(Q)|)$
@@ 4.1.1. Selection Strategy
FREE uses the dataset in the workload as the source of n-gram selection, and selects a prefix-free set of n-grams based on selectivity. Note that FREE does not use the query workload for n-gram selection. The candidate n-gram set $G$ is $G(D)=\bigcup_{d\in D}G(d)$. FREE decides whether an n-gram will be selected by the Usefulness criteria as defined below.
Definition 4.1 (FREE Usefulness).
By setting a fixed selectivity threshold $\mathbf{c}$, n-gram $g$ is deemed useful if its selectivity is less than the selectivity threshold, $c_{g}<\mathbf{c}$.
FREE makes two important assumptions for n-grams selection.
@@
FREE selects only n-grams that are useful. Since usefulness selection criteria is essentially a selectivity upper bound, if we know that an n-gram $g$ is useful, then any longer n-gram having $g$ as a substring is also useful. For example, if n-gram pdf is useful with a selectivity $c_{g}$, then n-grams such as $g^{\prime}=$ .pdf and $g^{\prime\prime}=$ pdfg will have selectivity $c(g^{\prime})<\mathbf{c}$ and $c(g^{\prime\prime})<\mathbf{c}$, therefore also useful.
@@
For a useful n-gram $g=g_{1}g_{2}\cdots g_{n_{1}}$ of size $n_{1}$ having selectivity $c_{g}$, any other longer n-grams of size $n_{2}$
@@
To effectively reduce the size of n-grams selected, FREE selects the prefix-minimal useful set of n-grams, deriving from the Apriori method of finding the maximal frequent sets in data mining literature (Agrawal et al., 1996). Essentially, it generates all candidate n-grams in increasing size order iteratively. For iteration $i$, it generates the candidate set of n-grams of length $i$ by extending all useless n-grams of length i1i-1 by one character, inserting the useful n-grams in the candidate set into the index, and using the useless ones in the candidate set for iteration $i+1$. This way, it is not necessary to generate all possible n-grams in each iteration. This method also ensures that the set of n-grams in the index is a prefix-minimal set, as the breadth-first search ensures that the shortest prefix of a useful n-gram is visited first. No n-grams in candidate sets of future iterations have the selected n-grams as prefixes, as the useful n-grams are never extended to generate a candidate n-gram.
FREE confines the search space by constraining the length $n$ of the n-grams selected. Briefly, there are two parameters to tune: 1) selectivity threshold $c$ that distinguish useful n-grams from useless n-grams, and 2) n-gram size $n$ to control the length of the index keys. In the original paper, the authors uses $c=0.1$ and $2\leq n\leq 10$ in their experiments. It also defined a simple query plan considering the index keys. We include a description of the query plan generation in Appendix A of the full paper. We implemented the original version of FREE with the query plan and an accompanied matcher in the codebase, but for a fair comparison of the n-gram selection methods, we do not use it for the experiments. The index data structure used is inverted index with n-grams as keys and posting lists as values.
@@ 4.1.2. Complexity Analysis
Let the candidate set at step $i$ be $M_{i}$. Initially, FREE scans the dataset to identify all unique unigrams and compute their selectivities, using a temporary hashmap of size $|M_{1}|=|\Sigma|$. This base step has time complexity $O(|D|+|\Sigma|)$. After filtering for useful unigrams, the remaining ones are extended to form bigrams $M_{2}$, with space and compute overheads of $O(|M_{2}|)$ and $O(|D|+|M_{2}|)$, respectively.
At each step $i$, the runtime space is $O(|M_{i}|)$ and compute time is $O(|D|+|M_{i}|)$. Since previous hashmaps are discarded, selecting n-grams up to size $k$ results in total space $O(|M_{k}|)$ and compute time $O(k\cdot(|D|+|M_{k}|))$. Given that $\sum_{i=1}^{k}|M_{i}|\leq|D|$, the overall space is $O(|D|)$ and compute time is $O(k\cdot|D|)$. In practice, $k$ is usually small and can be treated as a constant.
@@ 4.2.1. Selection Strategy
Definition 4.3 (BEST Cover).
For a workload $W=(Q,D)$, the cover of an n-gram $g$ is the set of data records $d\in D$ and query $q\in Q$ pairs defined as follows:
$cover(g)=\{(q,d)\in Q\times D\mid g\in q\land g\notin d\}$
For a set of n-grams $P$, the cover of the n-gram set is the union of the cover of each n-gram in the set.
$cover(P)=\bigcup_{g\in P}cover(g)$
The n-gram selection problem is then transformed to a budgeted maximum set cover problem (Khuller et al., 1999) that aims to find the set that maximize the cover over all sets $P$ that satisfy the budget constraint Formally, BEST define the value of indexing an n-gram with respective to the entire workload by its benefit.
Definition 4.4 (BEST Benefit).
For a workload $W=(Q,D)$ and an n-gram set $I$, the benefit of an n-gram $g\notin I$ is the number of additional query-data pairs covered by $g$ that is not already covered by $I$.
@@
To progressively calculate the benefit of each n-gram $g$ of $G(W)=G(D)\bigcup G(Q)$, we need two matrices to store the existence of $g$ in each data record $d$ and each query $q$. Specifically, the two matrices are of size $|Q|\cdot|G(W)|$ and $|D|\cdot|G(D)|$ respectively.
Definition 4.5 (BEST Cost).
For a workload $W$ consists of dataset $D$ and query set $Q$, the cost to index an n-gram $g$ is the storage overhead of $g$.
@@
Definition 4.6 (BEST Utility).
For a workload $W$ consists of dataset $D$ and query set $Q$ and an existing n-gram set $I$ on $W$, the utility of indexing an additional n-gram $g$ is the ratio of its benefit over cost, i.e., utility(g)=benefit(g,I)/cost(g)utility(g)={benefit(g,I)}/{cost(g)}
BEST selects n-grams based on their utility. The brute-force method would be to iteratively select the n-gram with the highest utility, $g_{max}$, based on the workload $W$ and the set of selected n-grams $I$, among all possible n-grams with a positive benefit.
@@
Therefore, according to the sparsity assumption Assumption BEST-3 and the definition of cover, instead of considering all n-grams in the query set and dataset, we can trim away the n-grams that only exists in the dataset. Therefore, the candidate n-gram set $G$ of BEST is $G(Q)$. BEST choose to use adjacency lists: $Q$-$G$-list and $G$-$D$-list rather than two matrices of sizes $|Q|\cdot|G(W)|$ and $|G(W)|\cdot|D|$ to store the existence of candidate n-grams to reduce space usage. The candidate n-grams, queries, and data records are each assigned an index number. The indices of $Q$-$G$-list correspond to query numbers, and each element is a list of n-gram numbers in this query. Similarly, the indices of $G$-$D$-list correspond to n-gram numbers, and each element is a list of data record numbers which the n-gram is in.